Titan Mining Corporation has 9.4 million shares of common stock outstanding and 380,000 4 percent semiannual bonds outstanding, par value $1,000 each. The common stock currently sells for $42 per share and has a beta of 1.2, and the bonds have 10 years to maturity and sell for 113 percent of par. The market risk premium is 8.2 percent, T-bills are yielding 3 percent, and the company tax rate is 35 percent.

a. What is the firms market value capital structure?

debt

equity
Answer #1

Equity Market value Weights $ 394,800,000 47.90% =9400000*42 Debt $ 429,400,000 52.10% =380000*1130 Total $ 824,200,000 100.0

*Please rate thumbs up